OpenStudy (anonymous):

what is the prime factorization of 18

OpenStudy (anonymous):

What are the factors of 18? That is, what numbers multiplied together yield 18?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Any two numbers that make up 18, that's what I am asking you :-)

OpenStudy (dcolley):

18 factors are:1,2,3,6,9,18

OpenStudy (dcolley):

1*18=18 2*9=18 3*6=18

OpenStudy (anonymous):

Ok, as @dcolley showed, those are the factors. Now, pick any one that has a prime factor: 2*9 or 3*6

OpenStudy (anonymous):

Then, we have to decomposo the non-prime factor. I will pick 3*6. Which two numbers multiplied together yield 6?

OpenStudy (anonymous):

decompose*

OpenStudy (dcolley):

I was thinking maybe 9 and 2 can you clarify why you chose 3*6

OpenStudy (anonymous):

 i have seen it done like this